OpenStudy (anonymous):

if 1/a +1/b=10, what is the value of a+b? a. 1/10 b. 2/5 c. 1 d. 10 e. cannot be determined

OpenStudy (raden):

1/a +1/b=10 1/a = 10 - 1/b 1/a = (10b - 1)/b a = b/(10b -1) if b = 1 get a = 1/9, then a+b = 1/9 + 1 = 10/9 if b = 2 get a = 2/19, then a+b = 2/19 + 2 = 40/19 ........ ........ i think it would has finitely solution here

OpenStudy (raden):

i means infinitely solutions, in other words cannot be determined

OpenStudy (whpalmer4):

Yeah, if you substitute that value of \(a\) into \(a+b\) you get \[\frac{b^2}{b-10}\]which has as many values as you want...